# Time-Series Foundation Model for Value-at-Risk

### v1

@@ 3.1 Experiment Setup
In this paper, we implement the recently introduced pre-trained time-series model, Google’s TimesFm, for forecasting value at risk and compare its performance against state-of-the-art parametric and non-parametric approaches, namely, GARCH model with the standardized residual modelled as empirical distribution function (G-EDF), normal distribution (G-N) and Student’s-t (G-t) distribution, one factor Generalized AutoRegressive Score model (GAS) and rolling window method (Historical).
@@
For each of the models, G-EDF, G-N, G-t and one factor GAS, we obtained the VaR estimates at levels 1%, 2.5%, 5% and 10% over the test period, 2nd January 2015 to 19th September 2023. For the rolling window (Historical) method, we used the data of 512 days preceding the first test day (i.e., 2nd January 2015), and forecasted the VaR at various levels. We then shift the training sample by one day forward in time to include more recent data and exclude the oldest data point such that a fixed size of the rolling window is maintained. At each rolling step, we obtain the prediction for the next day using the most recent 512 observations, thus resulting in a sequence of 2268 VaR estimates for each level. For GARCH and GAS models, we followed Patton et al., (2019) and used January 2005 to December 2014 as a training period for estimating model parameters, which were then used to forecast over the same test period as before, again resulting in a sequence of 2268 VaR estimates for each level.
@@
TimesFm is originally trained on a large number of public and synthetic datasets and exhibits robust out-of-the-box zero-shot performance in comparison to the accuracy of various state-of-the-art forecasting models specific to individual datasets under consideration. The model is flexible enough to forecast with different input length, prediction length and time granularities at inference time. The pre-trained model returns the point forecasts along with deciles, i.e., the quantilesAs per Google’e TimesFm github page, “We experimentally offer quantile heads but they have not been calibrated after pretraining.” at $10\%,20\%,\ldots,90\%$ levels. We experimented with different prediction horizons, but with a fixed input length of 2 years (512 days). Specifically, we set the prediction horizon to 1 day, 21 days and 63 days and denote these models as PT1, PT21 and PT63. For PT1, the input window shifts daily with the prediction points, while for PT21 and PT63, we use a fixed set of input observations for daily forecasts over the subsequent 21 and 63 days, respectively. For example, with PT21 (or PT63), we estimated VaR for a 21-day (or 63-day) period starting from January 2, 2015, using the most recent 512 observations. After this, we shifted the training data forward by 21 days (or 63 days) to incorporate more recent data and applied the pre-trained model for the next 21 (or 63) days, generating VaR estimates over the test period.
The variation in prediction horizon allows us to examine the model’s ability to generalize across forecasting horizons of varying lengths, i.e., short and long horizons. However, as mentioned above, the pre-trained models, PT1, PT21 and PT63, only produced VaR forecast at 90% level.
To better align the model with our objective, we conducted a fine-tuning procedure using the stock returns of S&P100 and its constituents. We divided the first period of ten years, January 2005 to December 2014, into two sets, first, the training set from January 2005 to December 2011 and second, the validation set from January 2012 to December 2014. We then fine-tuned the TimesFm model three times, each time starting from the pre-trained foundation model. More specifically, in each of the two sub-periods, training and validation samples of fixed context (input) length 512 and a test length of 1 were obtained. Then, the model was fine-tuned using the specifications described below. After getting the best fine-tuned model, we again experimented with different prediction horizons, but with a fixed input length of 2 years (512 days), each time starting from the pre-trained model. Therefore, similar to pre-trained model, we obtained three fine-tuned models, which we denote by FT1, FT21 and FT63. While finetuning, we modified the quantile levels to 0.01, 0.025, 0.05 and 0.1.
@@
We adopted the code “finetuning” in TimesFm library. More specifically, fine-tuning was carried out using a PatchedDecoderFinetuneModel with specific adjustments tailored for our task. During fine-tuning, we applied linear probing where only the core layer parameters were updated but the transformer layers were kept fixed. To prevent gradient explosion, we applied gradient clipping with a threshold of 100. The optimization process employed the Adam optimizer, that followed a cosine schedule for learning rate that started at $1\times 10^{-3}$ and decayed to $1\times 10^{-4}$ over 40,000 steps. Further, an exponential moving average decay of 0.9999 was also included in the optimization. The training loop ran for up to 100 epochs, with a provision of early stopping tied to the loss on the validation set and also a patience level was set to 5 epochs without improvement. Throughout the fine-tuning procedure, check-points were saved and finally the best model was restored for evaluation purposes.
@@ 4 Evaluation metrics
In financial risk management, VaR is a widely used measure that estimates the potential loss in the value of a stock over a given time period, with a certain confidence level. However, VaR estimates are erroraneous and hence often biased due to various sources of errors such as model errors, sampling errors, etc. Therefore, to ensure that VaR models are reliable, it is crucial to perform rigorous backtesting. Backtesting is the process of comparing the VaR predictions against actual observed outcomes to determine the accuracy of the model. Since a VaR value is fundamentally unobservable, we need to judge the performance of a model by cheking whether these estimates are consistent with the realized returns or not, given the confidence level at which these estimates were obtained. Thus, VaR backtesting is critical for evaluating the accuracy of the VaR forecasts.
@@
An alternative to AE ratio is the proportion of failure test (POF), also known as the unconditional coverage test (UC). which statistically evaluates whether the actual number of VaR violations matches the expected number of violations over time, based on the chosen confidence level. For instance, it calculates the failure rate defined as proportion of number of violations ($A$) to the total number of observation $T$ and examines it’s consistence with the models confidence level $\alpha$. More specifically, let $I_{t}$ denotes the indicator function that the realized return was less than the VaR estimate, i.e.,
$I_{t}=\begin{cases}1&\ r_{t}<q_{t}\\ 0,&\ \text{otherwise}\end{cases}$
where $r_{t}$ is the realized return at time $t$. Thus, $A=\sum_{t=1}^{T}I_{t}$ denotes the number of violations in the sample. From Kupiec et al., (1995), $A$ follows a binomial distribution, i.e., $A\sim B(T,\alpha)$, thus, under the null that $A/T=\alpha$, the relevant likelihood ratio statistic is
@@
As stated in Kupiec et al., (1995), the UC test is capable of rejecting a model for both low and high failures, but has poor power in general. Further, the UC test tests the number of violations, while it is cruicial to test whether these losses are spread evenly over time or not. The Conditional Coverage (CC) test, proposed by Christoffersen, (1998), does the same, that is, it tests whether the violations are randomly distributed over time and not clustered. That is, CC test takes into account any conditionality in VaR estimates: the model must respond to clustering events such as if volatilities are high in some period and low in others. The test statistics is given by
$-2\ln[(1-\alpha)^{T-A}\alpha^{A}]+2\ln[(1-\pi_{01})^{n_{00}}\pi_{01}^{n_{01}}(1-\pi_{11})^{n_{10}}\pi_{11}^{n_{11}}],$
@@
In the evaluation of VaR forecasts, the quantile score is a commonly used metric and is now standard in the VaR literature (Ehm et al., , 2016; Gneiting, , 2011; Taylor, , 2019). The above mentioned statistical tests are good to evaluate the statistical adequacy of VaR estimates, that is, a model for VaR forecasting is deemed to be an adequate model if we fail to reject the null hypothesis. However, these tests cannot conclude whether an ‘adequate’ model outperforms another ‘adequate’ model. Scoring functions allow us to do an honest assessment of VaR as a perfect VaR estimate would always minimize the scoring function. The quantile score $QS_{t}$ can be defined as follows:
@@
For every instant $t$, a model’s score increases either by (a) $\alpha-1$ fraction of the distance between the realized return and the VaR forecast, if realized return is smaller than the VaR forecast for the same period, or (b) by $\alpha$ fraction of the distance otherwise. This scoring rule is consistent for quantile estimation, meaning that it achieves its minimum value when the quantile forecast $q_{t}$ correctly estimates the true quantile of the distribution. Therefore, models that produce lower average quantile scores in the out-of-sample period are considered superior. Finally, to compare the performance of two models, we use Diebold–Mariano (DM) test (Diebold and Mariano, , 2002) with
@@ 6 Conclusion and Discussion
The question of whether foundation models can outperform traditional methods in VaR analysis is fundamentally important from the perspective of economics literature. In this paper, we evaluated the performance of Google’s foundation model for time series, TimesFM, in estimating Value-at-Risk (VaR) for the S&P 100 index and its 91 constituents. Our main finding is that in comparison to the state-of-the-art econometric models for VaR, TimesFM performs surprisingly well. Specifically, TimesFM outperforms all econometric benchmarks, including the GAS model, which is widely regarded as the leading method in recent econometric literature (Patton et al., , 2019), in terms of Actual vs. Expected violations (AE). Traditional models like GARCH (with different residual distributions) and empirical quantile estimates fall notably behind. These results remain consistent across different VaR levels, underscoring the robustness of TimesFM. Additionally, we assessed performance using the Quantile Score Loss Function, where TimesFM showed results comparable to the best econometric models.
We initially used the pre-trained TimesFM model without any task-specific training as a zero-shot predictor and observed significant statistical improvements after fine-tuning it with our data. The fine-tuned TimesFm model performed as well as the parametric, semi-parametric, and non-parametric approaches from the literature of forecasting the value at risk. As the results favor the foundation AI model, financial time-series modeling could shift toward data-driven approaches, partially reducing the need for mathematical modeling. Such a shift would have significant implications for both academia and the financial industry, as the application of foundation models can be applied with relatively minimal methodological expertise in econometrics. In conclusion, foundation models hold promise for improving Value-at-Risk (VaR) estimation, offering flexible alternatives to traditional econometric approaches.
@@
However, the adoption of foundation models is accompanied by notable challenges. A key challenge with these models is their “black-box” nature, which can make it difficult to understand how they arrive at their predictions. This is particularly concerning because we still have only a partial understanding of how these models operate, where they might fail, and what they are truly capable of, mainly due to their complex, emergent behaviors. This lack of transparency is especially problematic in high-stakes financial settings, where it’s crucial to understand the reasoning behind predictions for effective risk management and decision-making. Moreover, this opacity can create hurdles for regulatory acceptance, as regulators typically require clear and explainable models to meet compliance standards. Therefore, improving the interpretability of foundation models is critical to align with regulatory expectations and support their wider use in practice.
While pre-trained models like TimesFM can reduce initial training costs, fine-tuning these models still demands substantial computational resources. This requirement can limit access for smaller institutions or practitioners with limited computational capacity, presenting a barrier to widespread adoption. The resource-intensive nature of foundation models also raises concerns about scalability, as expanding their use could lead to increased costs and infrastructure demands. Finally, while time series foundation models show promise in adapting to out-of-distribution scenarios, similar to how humans use “common sense” reasoning, they are also prone to hallucinations and may produce decisions that appear reasonable but are incorrect (Chakraborty et al., , 2024). This highlights the importance of careful evaluation in real-world deployments.

### v2

@@ 3.1 Experiment Setup
We implement the recently introduced pre-trained time-series model, Google’s TimesFM, for forecasting Value at risk and compare its performance against state-of-the-art parametric and non-parametric approaches, namely, GARCH model with the standardized residual modelled as empirical distribution function (G-EDF), normal distribution (G-N) and Student’s-t (G-t) distribution, one factor Generalized AutoRegressive Score model (GAS) and rolling window method (Historical).
@@
For GARCH and GAS models, we followed Patton et al. (2019) and used January 2005 to December 2014 as a training period for estimating model parameters, which were then used to forecast over the test period from the 2nd January 2015 to the 19th September 2023, thus resulting in a sequence of 2,268 VaR estimates. For the rolling window (Historical) method, we used the data of 512 days preceding the first test day (i.e., 2nd January 2015), and forecasted the VaR at various levels. We then shift the training sample by one day forward in time to include more recent data and exclude the oldest data point such that a fixed size of the rolling window is maintained. At each rolling step, we obtain the prediction for the next day using the most recent 512 observations, thus resulting in a sequence of 2,268 VaR estimates for each level. For each of the models, G-EDF, G-N, G-t, one factor GAS, and Historical, we obtained the VaR estimates at levels 1%, 2.5%, 5%, and 10%.
@@
TimesFM is originally trained on a large number of public and synthetic datasets and exhibits robust out-of-the-box zero-shot performance in comparison to the accuracy of various state-of-the-art forecasting models specific to individual datasets under consideration. The model is flexible enough to forecast with different input length, prediction length and time granularities at inference time. The pre-trained model returns the point forecasts along with deciles, i.e., the quantiles [^6] As per Google’e TimesFM github page, “We experimentally offer quantile heads but they have not been calibrated after pretraining.” at $10\%,20\%,\ldots,90\%$ levels. We experimented with different prediction horizons with a fixed input length of 2 years (512 days). Specifically, we set the prediction horizon to 1 day, 21 days and 63 days and denote these models as PT1, PT21 and PT63. For PT1, the input window shifts daily with the prediction points, while for PT21 and PT63, we use a fixed set of input observations for daily forecasts over the subsequent 21 and 63 days, respectively. For example, with PT21 (or PT63), we estimated one-day VaRs over a 21-day (or 63-day) period starting from January 2, 2015, using the most recent 512 observations. After this, we shifted the training data forward by 21 days (or 63 days) to incorporate more recent data and applied the pre-trained model for the next 21 (or 63) days, generating one-day VaRs estimates over this period. The variation in prediction horizon allows us to examine the model’s ability to generalize across forecasting horizons of varying lengths, i.e., short and long horizons. However, as mentioned above, the pre-trained models, PT1, PT21 and PT63, only produced VaR forecast at 90% level.
@@
To better align the model with our objective, we conducted a fine-tuning procedure using the stock returns of S&P100 and its 91 constituents. We adopted the code “finetuning” in TimesFM library. More specifically, fine-tuning was carried out using a PatchedDecoderFinetuneModel with specific adjustments tailored for our task. During fine-tuning, we applied linear probing where only the core layer parameters were updated but the transformer layers were kept fixed. To prevent gradient explosion, we applied gradient clipping with a threshold of 100. The optimization process employed the Adam optimizer, that followed a cosine schedule for learning rate that started at $1\times 10^{-3}$ and decayed to $1\times 10^{-4}$ over 40,000 steps. Further, an exponential moving average decay of 0.9999 was also included in the optimization. The training loop ran for up to 100 epochs, with a provision of early stopping tied to the loss on the validation set and also a patience level was set to 5 epochs without improvement. Throughout the fine-tuning procedure, check-points were saved and finally the best model was restored for evaluation purposes.
In the fine-tuning process, we divided the first ten-year period, from January 2005 to December 2014, into two subsets: the training set from January 2005 to December 2011, and the validation set from January 2012 to December 2014. The validation data was used for hyperparameter optimization and to track the model’s performance after every epoch. That is, validation data is used to tune hyper-parameters by evaluating the performance on validation set during training, and selecting the best model check-point, that is, the one with lowest validation loss. With a fixed input length of 512, we fine-tuned the pre-trained model separately for three different prediction horizons. Specifically, we obtained the fine-tuned model FT1 using a prediction horizon of 1 day, FT21 with a horizon of 21 dats, and FT63 with a horizon of 63 days. Overall, three fine-tuned models were generated, denoted as FT1, FT21, and FT63, corresponding to the different prediction horizons. With fine-tuning we re-trained the model for the quantile levels 0.01, 0.025, 0.05, and 0.1 as the pre-trained models used quantiles 0.1, 0.2, … 0.9 only.
@@ 4 Evaluation metrics
In financial risk management, VaR is a widely used measure that estimates the potential loss in the value of a stock over a given time period, with a certain confidence level. However, VaR estimates are erroneous and hence often biased due to various sources of errors such as model errors, sampling errors, etc. Therefore, to ensure that VaR models are reliable, it is crucial to perform rigorous backtesting. Backtesting is the process of comparing the VaR predictions against actual observed outcomes to determine the accuracy of the model. Since a VaR value is fundamentally unobservable, we need to judge the performance of a model by checking whether these estimates are consistent with the realized returns or not, given the confidence level at which these estimates were obtained. Thus, VaR backtesting is critical for evaluating the accuracy of the VaR forecasts.
@@
An alternative to AE ratio is the proportion of failure test (POF), also known as the unconditional coverage test (UC). which statistically evaluates whether the actual number of VaR violations matches the expected number of violations over time, based on the chosen confidence level. For instance, it calculates the failure rate defined as proportion of number of violations ($A$) to the total number of observation $T$ and examines it is consistence with the models confidence level $\alpha$. More specifically, let $I_{t}$ denotes the indicator function that the realized return was less than the VaR estimate, i.e.,
$I_{t}=\begin{cases}1&\ r_{t}<q_{t}\\ 0,&\ \text{otherwise},\end{cases}$
where $r_{t}$ is the realized return at time $t$. Thus, $A=\sum_{t=1}^{T}I_{t}$ denotes the number of violations in the sample. From Kupiec (1995), $A$ follows a binomial distribution, i.e., $A\sim B(T,\alpha)$, thus, under the null that $A/T=\alpha$, the relevant likelihood ratio statistic is
@@
As stated in Kupiec (1995), the UC test is capable of rejecting a model for both low and high failures, but has poor power in general. Further, the UC test tests the number of violations, while it is crucial to test whether these losses are spread evenly over time or not. The Conditional Coverage (CC) test, proposed by Christoffersen (1998), does the same, that is, it tests whether the violations are randomly distributed over time and not clustered. That is, CC test takes into account any conditionality in VaR estimates: the model must respond to clustering events such as if volatilities are high in some period and low in others. The test statistics is given by
$-2\ln[(1-\alpha)^{T-A}\alpha^{A}]+2\ln[(1-\pi_{01})^{n_{00}}\pi_{01}^{n_{01}}(1-\pi_{11})^{n_{10}}\pi_{11}^{n_{11}}]$
@@
In the evaluation of VaR forecasts, the quantile score is a commonly used metric and is now standard in the VaR literature (Ehm et al., 2016; Gneiting, 2011; Taylor, 2019). The above mentioned statistical tests are good to evaluate the statistical adequacy of VaR estimates, that is, a model for VaR forecasting is deemed to be an adequate model if we fail to reject the null hypothesis. However, these tests cannot conclude whether an ‘adequate’ model outperforms another ‘adequate’ model. Scoring functions allow us to do an honest assessment of VaR as a perfect VaR estimate would always minimize the scoring function. The quantile score $QS_{t}$ can be defined as follows:
@@
For every instant $t$, a model’s score increases either by (a) $\alpha-1$ fraction of the distance between the realized return and the VaR forecast, if realized return is smaller than the VaR forecast for the same period, or (b) by $\alpha$ fraction of the distance otherwise. This scoring rule is consistent for quantile estimation, meaning that it achieves its minimum value when the quantile forecast $q_{t}$ correctly estimates the true quantile of the distribution. Therefore, models that produce lower average quantile scores in the out-of-sample period are considered superior. Finally, to compare the performance of two models, we use Diebold–Mariano (DM) test (Diebold and Mariano, 2002) with
@@ 6 Conclusion and discussion
The question of whether pre-trained foundation time-series models can outperform traditional econometric methodsis fundamentally important from the perspective of economics literature. In this paper, we evaluated the performance of Google’s foundation model for time series, TimesFM, in estimating Value-at-Risk (VaR) for the S&P 100 index and its 91 constituents. We initially used the pre-trained TimesFM model without any task-specific training as a zero-shot predictor and observed significant statistical improvements after fine-tuning it with our data. Our main finding is that in comparison to the state-of-the-art econometric models for VaR, fine-tuned TimesFM performs surprisingly well. It outperforms all econometric benchmarks, including the GAS model, which is widely regarded as the leading method in recent econometric literature (Patton et al., 2019), in terms of Actual vs. Expected violations (AE). Traditional models like GARCH (with different residual distributions) and empirical quantile estimates fall notably behind. These results remain consistent across different VaR levels, underscoring the robustness of TimesFM. Additionally, we assessed performance using the Quantile Score Loss Function, where TimesFM showed results comparable to the best econometric models.
@@
As the results favor the foundation AI model, financial time-series modeling could shift toward data-driven approaches, partially reducing the need for mathematical modeling. Such a shift would have significant implications for both academia and the financial industry, as the application of foundation models can be applied with relatively minimal methodological expertise in econometrics. However, the adoption of foundation models is accompanied by notable challenges. A key challenge with these models is their “black-box” nature, which can make it difficult to understand how they arrive at their predictions. This is particularly concerning because we still have only a partial understanding of how these models operate, where they might fail, and what they are truly capable of, mainly due to their complex, emergent behaviors. While time series foundation models show promise in adapting to out-of-distribution scenarios, similar to how humans use “common sense” reasoning, they are also prone to hallucinations and may produce decisions that appear reasonable but are incorrect (Chakraborty et al., 2024). This lack of transparency is especially problematic in high-stakes financial settings, where it is crucial to understand the reasoning behind predictions for effective risk management and decision-making. Moreover, it can create hurdles for regulatory acceptance, as regulators typically require clear and explainable models to meet compliance standards. Therefore, improving the interpretability of foundation models is critical to align with regulatory expectations and support their wider use in practice.
Overall, our results—which are highly generalizable, having been conducted across over 90 assets with multiple settings—favor the pre-trained foundation model, highlighting its potential to enhance Value-at-Risk (VaR) estimation by offering a completely new alternative to traditional econometric methods. We anticipate further research expanding into various econometric and financial challenges, such as portfolio optimization and volatility forecasting, but studies that examine the risks associated with the use of foundation models on econometric problems closely and critically.